[ If g(x) = 3.5x^3
find: g(4) ]

Found 2 solutions by richwmiller, TimothyLamb:
Answer by richwmiller(17219)

plug in 4 for x
g(4)=3.5*4^3
g(4)=7/2*64
g(4)=7*32

Answer by TimothyLamb(4379)

g(x) = 3.5x^3
---
g(4) = 3.5(4)^3
g(4) = 3.5(64)
---
g(4) = 224
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php